Jeff takes 5 hr longer to build a fence than it takes Bill. When they work together, it takes them 6 hours. How long would it take Bill to do the job alone?

Answer by checkley77(12844)

X(X+5)/(X+X+5)=6
(X^2+5X)/(2X+5)=6
X^2+5X=6(2X+5)
X^2+5X=12X+30
X^2+5X-12X-30=0
X^2-7X-30=0
(X-10)(X+3)=0
X-10=0
X=10 HOURS FOR BILL TO COMPLETE THE JOB ALONE.
PROOF:
10*15/(10+15)=6
150/25=6
6=6